Consider a person sitting nude on a beach in Florida. On a sunny day, visible radiation energy from the sun is absorbed by the person at a rate of 30 kcal/h or 34.9 W. The air temperature is a warm 30 ℃ and the individual’s skin temperature is 32 ℃. The effective body surface exposed to the sun is 0.9
1
Expert's answer
2020-11-25T10:46:31-0500

Since no question was asked, let us find the net energy gain from solar thermal radiation each hour:


"q=\frac{P}{A}=\frac{34.9}{0
.9}=38.8\text{ W/m}^2,\text{ or J/(s}\cdot\text{m)}^2"

The energy gain each hour will be


"q=38.8\cdot60=2328\text{ J/(m}^2\cdot\text{h})"